2) A concessionaire for the local ballpark has developed a table of conditional values for the various alternatives (stocking decision) and states of nature (size of crowd). All values are in US Dollars. LargeAverage Small 3 Probability Alternative 1 9,000 12,0002,000 Aisanativ:l» t,0008,000 &000 Alternative 3 5,00012,0005,000 If the probabilities associated with the states of nature are 0.30 for a large crowd, 0.50 for an average crowd, and 0.20 for a small crowd, determine: (a) the alternative that provides the greatest expected monetary value (EMV). (b) the expected value of perfect information (EVPI). 4
Answer #1

(a)

EMV of alternative 1 = 0.3*9000+0.5*12000+0.2*(-2000) = 8,300

EMV of alternative 2 = 0.3*22000+0.5*6000+0.2*6000 = 10,800

EMV of alternative 3 = 0.3*15000+0.5*12000+0.2*5000 = 11,500

Alternative 3 provides the greatest EMV (11,500)

________________________________________

(b) EVwPI = 0.3*MAX(9000,22000,15000)+0.5*MAX(12000,6000,12000)+0.2*MAX(-2000,6000,5000)

= 13,800

EVPI = EVwPI - Max EMV

= 13800 - 11500

= 2,300